From frequency distribution table given below, find-
i) mean
ii) median
iii) mode
iv) variance
v) standard deviation
L1 — L2 50–52 53–55 56–58 59–61 62–64
f 5 10 21 8 6
1
Expert's answer
2019-04-25T09:15:26-0400


(i):


"L_x =(L_2+L_1)\/2""f_x=f \cdot L_x""Mean= \frac {\sum f_x} {\sum f}=2850/50=57"

(ii):


"\text {F - cumulative frequency}""\frac {\sum f} {2} = 25 \rightarrow \text { class median: 56-58}""L_m=55.5 \text { - the lower boundary of the class median}""f_m=21 \text { - the frequency of the class median}""F_m=15 \text { - the cumulative frequency before class median}""w=3 \text { - the class width}""Median=L_m + \frac {\frac {\sum f} {2} - F_m} {f_m} w = 55.5 + \frac {25-15} {21} 3 = 56.929"

(iii):


"f_{max} = 21 \rightarrow \text { modal group: 56-58}""L_{md}=55.5 \text { - the lower boundary of the modal group}""f_{md} = 21 \text { - the frequency of the modal group}""Mode = L_{md} + \frac {f_{md} - f_{md-1}} {(f_{md} - f_{md-1})+(f_{md} - f_{md+1})} w=""=55.5+\frac {21-10} {21-10+21-8} 3=56.875"

(iv):


"Variance=\sigma ^2 = \frac {\sum f_x ^2} {\sum f} - (\frac {\sum f_x } {\sum f})^2 = \frac {2162718} {50} - 57^2 = 40005.36"

(v):


"\text {Standard deviation} = \sqrt{\sigma^2}=200.013"